# A Study on the Calibration of In-context Learning

### v3

@@ 4.4 Ablation Studies
Ablation with model sizes. As we enlarge the size of models, they become more confident and accurate (Fig. 2). As a result, the entropy decreases and ECE increases, showing that token-level calibration might have an inverse scaling relationship with model sizes.
Table 5: Qualitative Results of LLaMA on Commonsense and OpenBook
Dataset Case
Reasoning
$\hat{y}$ $p$
@@
7B
If you have curiosity but are blind and paralyzed, you would go somewhere to see things. You cannot hear news, analyse, go somewhere, examine thing or see favorite show.
C 0.5046
13B
If you are blind and paralyzed, you can’t see anything. So, you can’t go somewhere. You can’t examine thing. You can’t see favorite show. You can’t analyse. So, you can hear news.
A 0.8684
30B
If you are blind and paralyzed, you cannot see your favorite show. You cannot hear news as you are blind. You cannot examine thing as you are paralyzed. You can go somewhere to see things.
C 0.9720
@@
7B
A girl and her mom have the same hair length.
D 0.6365
13B
A girl and her mom have the same date of birth.
A 0.9782
30B
A girl and her mom have the same genes.
A 0.9831
Examine the hidden state and confidence score. To better understand the miscalibration issue of ICL, we conduct fined-grained experiments to examine ICL properties: we measure the norm of the representation vectors [^1] The representation vector refers to the intermediate output before the linear prediction layer. for different number of shots in ICL. Meanwhile, we also measure the confidence and entropy of the prediction for $y_{n}$, and the results are summarized in Tab. 4. When switching from 0-shot to 1-shot, all three measurements (representation norm, entropy, and confidence) drastically change; on the other hand, when $k$ increases from $1\rightarrow 4\rightarrow 8$, the change of measures would become smoother. Our discovery shows that adding in-context example can substantially impact model behaviors while model behaves relatively similar for various shots once the task is specified ($k\neq 0)$. Meanwhile, more ICL samples lead to smaller entropy and higher confidence in most cases.
Confidence and wrongly classified reasoning examples. To inspect the failure modes of LMs, we randomly sample 100 reasoning examples of LLaMA and plot the distribution of wrongly predicted samples and the confidence scores via thresholding. Similar to previous observations, as model sizes and the number of ICL examples scale up, LMs would generate more confident samples (Fig. 4 (c, d)). We observe “inverse scaling” behaviors where models with larger sizes are more error-prone and tend to generate more confidently wrong samples (Fig. 5).
Examples of hallucinated explanations for highly confident predictions. Next, we showcase in Tab. 5 that models generate both wrong explanations and incorrect predictions with high confidence. We also observe that most of the wrong predictions are highly confident, thus we manually examine the correctness of explanations on commonsense QA, and found its high correlations with predicted answer accuracy, which is the opposite of token-level explainability that tends to get worse when the accuracy improves.
@@ 5 Discussion and Concluding Remarks
Understanding the internal mechanism of ICL for calibration.
We observe that existing scaling recalibration methods cannot fully resolve the miscalibration issues of ICL, so better understanding and mitigation strategies are needed. A potential approach can be leveraging transparency tools and studying whether predictable errors exist during text generation.
@@ Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for their valuable feedback. Kempner Institute computing resources enabled this work.
Appendix A Additional Experimental Details
We provide prompts we adopt for experiments in Tab.6. Additional reliability plots are shown in Fig. 6.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ Appendix A Additional Experimental Details
SST-2
Review: it may not be a great piece of filmmaking, but its power comes from its soul’s - eye view of how well-meaning patronizing masked a social injustice, at least as represented by this case .
@@
CB
A: No, not really. I spend a lot of time with our income tax, though. especially, this year and last year. Um, I have been married for just a few years, so I’ve had to really switch around from the EZ form to the, uh, B: Schedule A. A: Right. B: Well, yeah. A: All the deductions and all that. B: Did you notice that when they passed the new simplified tax act, it seemed like it made everything harder?
@@
RTE
The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
@@
Strategy QA
Question: Can spiders help eggplant farmers control parasites? Choose the answer from True and False.
@@
Commonsense QA
"Question: Dan was a farmer with just one heifer. But that was okay, he only kept her for milk, and he didn’t think he’d find good farmland in a place as cold as where?

### v4

@@ 4.4 Ablation Studies
Ablation with model sizes. As we enlarge the size of models, they become more confident (as measured by the confidence histogram) and accurate (Fig. 2). Moreover, the ECE first increases and then decreases. In some settings like SST-2 and OpenBookQA, calibration errors may have a negative correlation with model sizes (Appendix Tab.9).
Dataset Case Reasoning $\hat{y}$ $p$
@@
7B If you have curiosity but are blind and paralyzed, you would go somewhere to see things. You cannot hear news, analyse, go somewhere, examine thing or see favorite show. C 0.5046
13B If you are blind and paralyzed, you can’t see anything. So, you can’t go somewhere. You can’t examine thing. You can’t see favorite show. You can’t analyse. So, you can hear news. A 0.8684
30B If you are blind and paralyzed, you cannot see your favorite show. You cannot hear news as you are blind. You cannot examine thing as you are paralyzed. You can go somewhere to see things. C 0.9720
@@
7B A girl and her mom have the same hair length. D 0.6365
13B A girl and her mom have the same date of birth. A 0.9782
30B A girl and her mom have the same genes. A 0.9831
Table 6: Qualitative Results of LLaMA on Commonsense and OpenBook
Examine the hidden state and confidence score. To better understand the miscalibration issue of ICL, we conduct fined-grained experiments to examine ICL properties: we measure the norm of the representation vectors [^1] The representation vector refers to the intermediate representation before the linear prediction layer. for different numbers of shots in ICL. Meanwhile, we also measure the confidence and entropy of the prediction for $y_{n}$, and the results are summarized in Tab. 5. When switching from 0-shot to 1-shot, all three measurements (representation norm, entropy, and confidence) drastically change; on the other hand, when $k$ increases ($1\rightarrow 4\rightarrow 8$), the change of measures would become smoother. Our discovery shows that adding in-context examples can substantially impact model behaviors while the model behaves relatively similarly for various shots once the task is specified ($k\neq 0)$. Meanwhile, more ICL samples lead to smaller entropy and higher confidence in most cases. Considering the alterations in feature representation, which can manifest in either an augmentation of the representation’s norm or a shift in direction, quantifying changes in feature direction poses challenges. Thus, we have chosen to examine changes in the norm as a surrogate measure, suggesting that as the number of ICL samples increases, there is a systematic alteration in the model’s features.
Confidence and wrongly classified reasoning examples. To inspect the failure modes of LMs, we randomly sample 100 reasoning examples of LLaMA and plot the distribution of wrongly predicted samples and the confidence scores via thresholding. Similar to previous observations, as model sizes and the number of ICL examples scale up, LMs would generate more confident samples (Fig. 4 (c, d)). We observe behaviors where models with larger sizes may be more error-prone and tend to generate more confidently wrong explanatory samples (Fig. 5).
Examples of hallucinated explanations for highly confident predictions. Next, we showcase in Tab. 6 that models generate both wrong explanations and incorrect predictions with high confidence. We also observe that most of the wrong predictions are highly confident, thus we manually examine the correctness of explanations on commonsense QA, and found its high correlations with predicted answer accuracy, which is the opposite of token-level explainability that tends to get worse when the accuracy improves. For additional qualitative examination of LLaMA’s performance on Strategy QA and WorldTree, please refer to Table 12.
@@ 5 Discussion and Concluding Remarks
Limitations. We acknowledge the need to expand our evaluation, which is primarily focused on QA and classification tasks, beyond existing open-sourced language models and datasets. Moreover, we didn’t consider nuances such as inherent disagreement about labels (Baan et al., 2022) and adaptive calibration error measures (Nixon et al., 2019) that might be important in certain use cases: it’s worth noting that situations may arise where multiple labels share the highest predicted probability. In such instances, the definition (Eq. (2)) doesn’t automatically become false; instead, we opt for the first maximal probability. These cases are less likely to occur in most of our experimental setups, where a substantial margin consistently exists between different labels.
@@ Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for their valuable feedback. HZ is supported by an Eric and Susan Dunn Graduate Fellowship. YY acknowledges support from the joint Simons Foundation-NSF DMS grant #2031899. SK acknowledges support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. Kempner Institute computing resources enabled this work. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
Appendix A Extended Related Work
Appendix B Additional Experimental Details
We provide prompts we adopt for experiments in Tab.7. Additional reliability plots are shown in Fig. 6. Moreover, we provide extra results that extend those in the main text. Our implementation is open-sourced at https://github.com/hlzhang109/icl-calibration. The greatest accuracy and ECE values are highlighted in bold and red, respectively. Extremely poor performance due to length truncation is omitted.
Model performance and calibration. We present experimental results considering different model sizes for text classification and reasoning in Tables 8 and 9, respectively. With the increase in model sizes, we observed overall improvements in model performance across most datasets. However, the calibration error (ECE) did not decrease immediately: for low-shot settings where $k<4$, models tend to have an ECE larger than $0.1$. On the other hand, ECE can decrease given more ICL examples ($k=8$) if context length is adequate. Overall, zero-shot ICL can lead to good calibration results though the predictive performance is substantially weaker. Interestingly, for some benchmarks like SST-2 and OpenBook QA, the ECE of the 30B model even surpassed that of the 7B model. Moreover, the ECE curves of the 7B and 13B models exhibited similar patterns to the 30B results as the number of ICL samples increased, as shown in the main Tab. (1).
The effect of fine-tuning. We provide full results of all finetuned LLMs in Table 11, complementing Fig. (3). We reach a similar conclusion as we explain in the main text: with an increasing number of ICL examples, accuracy generally improves but ECE first increases then decreases and miscalibration is widespread; an MoE model can also have the same accuracy-calibration trade-off; fine-tuning substantially improves accuracy but hurts calibration by a large margin.
Results reliability. Furthermore, as prompting is susceptible to various forms of biases and noises (Zhao et al., 2021; Han et al., 2023; Fei et al., 2023; Zhou et al., 2023a), to provide a comprehensive understanding of the experimental outcomes, we delve into the variance across all experimental repetitions. Table 10 provides a detailed analysis of the variance metrics, affirming the stability and reliability of our experimental findings.
@@ Appendix B Additional Experimental Details
SST-2 Review: it may not be a great piece of filmmaking, but its power comes from its soul’s - eye view of how well-meaning patronizing masked a social injustice, at least as represented by this case .
@@
CB A: No, not really. I spend a lot of time with our income tax, though. especially, this year and last year. Um, I have been married for just a few years, so I’ve had to really switch around from the EZ form to the, uh, B: Schedule A. A: Right. B: Well, yeah. A: All the deductions and all that. B: Did you notice that when they passed the new simplified tax act, it seemed like it made everything harder?
@@
RTE The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
@@
Strategy QA Question: Can spiders help eggplant farmers control parasites? Choose the answer from True and False.
@@
Commonsense QA "Question: Dan was a farmer with just one heifer. But that was okay, he only kept her for milk, and he didn’t think he’d find good farmland in a place as cold as where?
@@
Metric Dataset Model Size 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot
ECE AGNews 7B 0.067 0.105 0.225 0.158 0.086 0.075
13B 0.093 0.084 0.069 0.121 0.103 0.045
30B 0.261 0.043 0.049 0.067 0.049 0.047
CB 7B 0.133 0.218 0.172 0.197 0.202 0.215
13B 0.029 0.257 0.282 0.221 0.263 0.216
30B 0.069 0.312 0.216 0.217 0.192 0.181
RTE 7B 0.068 0.075 0.098 0.112 0.091 0.064
13B 0.042 0.104 0.048 0.048 0.049 0.050
30B 0.023 0.051 0.060 0.050 0.048 0.058
SST-2 7B 0.038 0.142 0.132 0.121 0.108 0.064
13B 0.051 0.134 0.108 0.084 0.073 0.053
30B 0.083 0.163 0.139 0.126 0.112 0.080
ACC AGNews 7B 0.447 0.629 0.563 0.630 0.777 0.833
13B 0.490 0.812 0.773 0.720 0.775 0.847
30B 0.370 0.830 0.817 0.810 0.821 0.855
CB 7B 0.482 0.596 0.675 0.696 0.691 0.729
13B 0.554 0.627 0.659 0.691 0.611 0.709
30B 0.500 0.696 0.789 0.834 0.814 0.796
RTE 7B 0.552 0.668 0.653 0.646 0.653 0.698
13B 0.679 0.673 0.708 0.723 0.723 0.746
30B 0.672 0.742 0.747 0.738 0.748 0.752
SST-2 7B 0.483 0.799 0.877 0.908 0.917 0.954
13B 0.483 0.918 0.943 0.955 0.962 0.969
30B 0.607 0.930 0.940 0.961 0.964 0.964
Table 8: Accuracy and Calibration of LLaMA model with three sizes across four text classification datasets.
Metric Dataset Model Size 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot
ECE Commonsense QA 7B 0.070 0.155 0.237 0.227 0.238 -
13B 0.066 0.161 0.282 0.292 0.310 -
30B 0.048 0.232 0.290 0.253 0.283 -
OpenBook QA 7B 0.040 0.241 0.270 0.184 0.130 0.121
13B 0.031 0.132 0.217 0.209 0.191 0.175
30B 0.048 0.232 0.290 0.253 0.283 -
Strategy QA 7B 0.133 0.275 0.206 0.243 0.242 0.227
13B 0.051 0.154 0.170 0.192 0.188 0.190
30B 0.204 0.154 0.174 0.172 0.161 0.193
World Tree 13B 0.065 0.113 0.226 0.250 0.284 -
30B 0.112 0.211 0.251 0.185 0.206 -
7B 0.074 0.124 0.198 0.179 0.203 -
ACC Commonsense QA 7B 0.224 0.292 0.388 0.421 0.406 -
13B 0.320 0.478 0.549 0.574 0.562 -
30B 0.356 0.589 0.608 0.675 0.644 -
OpenBook QA 7B 0.308 0.298 0.376 0.417 0.454 0.480
13B 0.362 0.454 0.509 0.551 0.580 0.611
30B 0.386 0.561 0.604 0.644 0.648 0.662
Strategy QA 7B 0.566 0.488 0.554 0.550 0.562 0.575
13B 0.554 0.598 0.621 0.595 0.618 0.612
30B 0.450 0.619 0.654 0.660 0.672 0.662
World Tree 7B 0.302 0.298 0.326 0.384 0.362 -
13B 0.444 0.437 0.495 0.519 0.492 -
30B 0.534 0.570 0.621 0.680 0.646 -
Table 9: Accuracy and Calibration of LLaMA models across three sizes on four reasoning datasets.
Dataset Metric 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot
CB ACC $0.500_{\pm 0.000}$ $0.696_{\pm 0.304}$ $0.789_{\pm 0.138}$ $\mathbf{0.834_{\pm 0.068}}$ $0.814_{\pm 0.068}$ $0.796_{\pm 0.110}$
ECE $0.143_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.409_{\pm 0.041}}$ $0.216_{\pm 0.061}$ $0.217_{\pm 0.057}$ $0.376_{\pm 0.053}$ $0.359_{\pm 0.071}$
RTE ACC $0.672_{\pm 0.000}$ $0.742_{\pm 0.018}$ $0.747_{\pm 0.032}$ $0.738_{\pm 0.044}$ $0.748_{\pm 0.043}$ $\mathbf{0.752_{\pm 0.039}}$
ECE $0.023_{\pm 0.000}$ $0.051_{\pm 0.020}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.060_{\pm 0.021}}$ $0.050_{\pm 0.023}$ $0.048_{\pm 0.017}$ $0.058_{\pm 0.022}$
SST-2 ACC $0.607_{\pm 0.000}$ $0.930_{\pm 0.025}$ $0.940_{\pm 0.066}$ $0.961_{\pm 0.017}$ $\mathbf{0.964_{\pm 0.012}}$ $\mathbf{0.964_{\pm 0.011}}$
ECE $0.106_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.339_{\pm 0.026}}$ $0.139_{\pm 0.058}$ $0.126_{\pm 0.053}$ $0.310_{\pm 0.022}$ $0.287_{\pm 0.014}$
AGnews ACC $0.370_{\pm 0.000}$ $0.830_{\pm 0.015}$ $0.817_{\pm 0.017}$ $0.810_{\pm 0.056}$ $0.821_{\pm 0.029}$ $\mathbf{0.855_{\pm 0.017}}$
ECE $0.261_{\pm 0.000}$ $0.043_{\pm 0.009}$ $0.049_{\pm 0.016}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.067_{\pm 0.029}}$ $0.049_{\pm 0.017}$ $0.047_{\pm 0.018}$
OpenBook QA ACC $0.386_{\pm 0.000}$ $0.561_{\pm 0.028}$ $0.604_{\pm 0.027}$ $0.644_{\pm 0.016}$ $0.648_{\pm 0.018}$ $\mathbf{0.662_{\pm 0.031}}$
ECE $0.036_{\pm 0.000}$ $0.231_{\pm 0.049}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.255_{\pm 0.050}}$ $0.207_{\pm 0.041}$ $0.206_{\pm 0.019}$ $0.191_{\pm 0.022}$
CommonSense QA ACC $0.356_{\pm 0.000}$ $0.586_{\pm 0.028}$ $0.608_{\pm 0.013}$ $\mathbf{0.675_{\pm 0.027}}$ $0.644_{\pm 0.034}$ $0.653_{\pm 0.090}$
ECE $0.048_{\pm 0.000}$ $0.232_{\pm 0.102}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.290_{\pm 0.022}}$ $0.253_{\pm 0.028}$ $0.283_{\pm 0.045}$ $0.289_{\pm 0.140}$
Strategy QA ACC $0.450_{\pm 0.000}$ $0.619_{\pm 0.030}$ $0.654_{\pm 0.033}$ $0.660_{\pm 0.022}$ $\mathbf{0.672_{\pm 0.015}}$
ECE $0.204_{\pm 0.000}$ $0.154_{\pm 0.029}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.174_{\pm 0.070}}$ $0.172_{\pm 0.025}$ $0.161_{\pm 0.008}$
World Tree ACC $0.554_{\pm 0.000}$ $0.570_{\pm 0.056}$ $0.621_{\pm 0.109}$ $\mathbf{0.680_{\pm 0.072}}$ $0.504_{\pm 0.074}$
ECE $0.112_{\pm 0.000}$ $0.211_{\pm 0.042}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.251_{\pm 0.101}}$ $0.185_{\pm 0.048}$ $0.144_{\pm 0.051}$
Table 10: The full results (mean and standard deviation) for various experimental configurations extending Table. 1.
Dataset Metric 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot
CB
Alpaca-7B ACC $0.552_{\pm 0.000}$ $0.668_{\pm 0.032}$ $0.653_{\pm 0.079}$ $0.646_{\pm 0.086}$ $0.653_{\pm 0.067}$ $\mathbf{0.698_{\pm 0.028}}$
ECE $0.016_{\pm 0.000}$ $0.119_{\pm 0.018}$ $0.123_{\pm 0.044}$ $0.122_{\pm 0.031}$ $0.115_{\pm 0.017}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.127_{\pm 0.020}}$
LLama2-Chat-7B ACC $0.375_{\pm 0.000}$ $0.566_{\pm 0.129}$ $0.643_{\pm 0.107}$ $0.670_{\pm 0.126}$ $\mathbf{0.677_{\pm 0.113}}$ $\mathbf{0.677_{\pm 0.111}}$
ECE $0.287_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.223_{\pm 0.078}}$ $0.170_{\pm 0.062}$ $0.153_{\pm 0.054}$ $0.154_{\pm 0.054}$ $0.170_{\pm 0.054}$
LLama2-7B ACC $0.339_{\pm 0.000}$ $0.464_{\pm 0.193}$ $0.511_{\pm 0.163}$ $0.538_{\pm 0.113}$ $0.534_{\pm 0.109}$ $\mathbf{0.575_{\pm 0.059}}$
ECE $0.125_{\pm 0.000}$ $0.222_{\pm 0.190}$ $0.174_{\pm 0.029}$ $0.206_{\pm 0.066}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.226_{\pm 0.071}}$ $0.222_{\pm 0.058}$
Mistral-7B-v0.1 ACC $0.500_{\pm 0.000}$ $0.643_{\pm 0.264}$ $0.725_{\pm 0.198}$ $\mathbf{0.827_{\pm 0.067}}$ $0.793_{\pm 0.063}$ $0.793_{\pm 0.121}$
ECE $0.063_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.330_{\pm 0.118}}$ $0.228_{\pm 0.094}$ $0.244_{\pm 0.036}$ $0.193_{\pm 0.048}$ $0.144_{\pm 0.028}$
vicuna-7b-v1.5 ACC $0.571_{\pm 0.000}$ $0.668_{\pm 0.049}$ $0.663_{\pm 0.052}$ $0.668_{\pm 0.058}$ $\mathbf{0.675_{\pm 0.061}}$ $0.648_{\pm 0.073}$
ECE $0.051_{\pm 0.000}$ $0.176_{\pm 0.034}$ $0.172_{\pm 0.047}$ $0.169_{\pm 0.054}$ $0.170_{\pm 0.047}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.181_{\pm 0.052}}$
AGNews
Alpaca-7B ACC $0.810_{\pm 0.000}$ $0.793_{\pm 0.041}$ $0.710_{\pm 0.110}$ $0.715_{\pm 0.111}$ $0.782_{\pm 0.079}$ $\mathbf{0.832_{\pm 0.029}}$
ECE $0.043_{\pm 0.000}$ $0.123_{\pm 0.033}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.190_{\pm 0.095}}$ $0.167_{\pm 0.093}$ $0.112_{\pm 0.057}$ $0.065_{\pm 0.019}$
LLama2-Chat-7B ACC $0.793_{\pm 0.000}$ $0.809_{\pm 0.031}$ $0.823_{\pm 0.046}$ $0.829_{\pm 0.035}$ $0.829_{\pm 0.028}$ $\mathbf{0.843_{\pm 0.019}}$
ECE $0.164_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.162_{\pm 0.030}}$ $0.143_{\pm 0.039}$ $0.138_{\pm 0.033}$ $0.138_{\pm 0.024}$ $0.127_{\pm 0.013}$
LLama2-7B ACC $0.573_{\pm 0.000}$ $0.832_{\pm 0.022}$ $0.789_{\pm 0.112}$ $0.801_{\pm 0.108}$ $0.849_{\pm 0.057}$ $\mathbf{0.868_{\pm 0.009}}$
ECE $0.102_{\pm 0.000}$ $0.037_{\pm 0.012}$ $0.074_{\pm 0.083}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.078_{\pm 0.082}}$ $0.052_{\pm 0.024}$ $0.053_{\pm 0.011}$
Mistral-7B-v0.1 ACC $0.780_{\pm 0.000}$ $0.847_{\pm 0.017}$ $0.842_{\pm 0.028}$ $0.820_{\pm 0.056}$ $0.808_{\pm 0.085}$ $\mathbf{0.867_{\pm 0.004}}$
ECE $0.193_{\pm 0.000}$ $0.059_{\pm 0.012}$ $0.044_{\pm 0.010}$ $0.052_{\pm 0.022}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.077_{\pm 0.049}}$ $0.043_{\pm 0.010}$
vicuna-7b-v1.5 ACC $0.740_{\pm 0.000}$ $0.803_{\pm 0.013}$ $0.834_{\pm 0.031}$ $0.824_{\pm 0.054}$ $\mathbf{0.835_{\pm 0.030}}$ $0.832_{\pm 0.036}$
ECE $0.063_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.139_{\pm 0.012}}$ $0.108_{\pm 0.025}$ $0.116_{\pm 0.034}$ $0.114_{\pm 0.014}$ $0.109_{\pm 0.034}$
RTE
Alpaca-7B ACC $0.672_{\pm 0.000}$ $0.644_{\pm 0.015}$ $0.687_{\pm 0.019}$ $0.696_{\pm 0.020}$ $\mathbf{0.703_{\pm 0.015}}$ $0.690_{\pm 0.035}$
ECE $0.175_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.270_{\pm 0.018}}$ $0.212_{\pm 0.034}$ $0.197_{\pm 0.028}$ $0.184_{\pm 0.026}$ $0.193_{\pm 0.025}$
LLama2-Chat-7B ACC $0.729_{\pm 0.000}$ $0.685_{\pm 0.042}$ $0.687_{\pm 0.048}$ $0.699_{\pm 0.040}$ $0.709_{\pm 0.034}$ $\mathbf{0.731_{\pm 0.033}}$
ECE $0.165_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.218_{\pm 0.031}}$ $0.205_{\pm 0.033}$ $0.198_{\pm 0.033}$ $0.184_{\pm 0.030}$ $0.172_{\pm 0.020}$
LLama2-7B ACC $0.682_{\pm 0.000}$ $0.684_{\pm 0.034}$ $\mathbf{0.698_{\pm 0.049}}$ $0.676_{\pm 0.058}$ $0.689_{\pm 0.068}$ $0.685_{\pm 0.050}$
ECE $0.044_{\pm 0.000}$ $0.076_{\pm 0.021}$ $0.084_{\pm 0.029}$ $0.085_{\pm 0.034}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.105_{\pm 0.031}}$ $0.083_{\pm 0.032}$
Mistral-7B-v0.1 ACC $0.686_{\pm 0.000}$ $0.731_{\pm 0.025}$ $0.756_{\pm 0.015}$ $0.768_{\pm 0.019}$ $\mathbf{0.776_{\pm 0.016}}$ $0.773_{\pm 0.025}$
ECE $0.054_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.121_{\pm 0.047}}$ $0.080_{\pm 0.042}$ $0.084_{\pm 0.033}$ $0.087_{\pm 0.025}$ $0.085_{\pm 0.035}$
vicuna-7b-v1.5 ACC $0.610_{\pm 0.000}$ $0.731_{\pm 0.015}$ $0.756_{\pm 0.011}$ $0.762_{\pm 0.013}$ $0.765_{\pm 0.019}$ $\mathbf{0.770_{\pm 0.026}}$
ECE $0.234_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.101_{\pm 0.021}}$ $0.073_{\pm 0.028}$ $0.067_{\pm 0.016}$ $0.057_{\pm 0.015}$ $0.052_{\pm 0.015}$
SST-2
Alpaca-7B ACC $0.730_{\pm 0.000}$ $0.868_{\pm 0.088}$ $0.939_{\pm 0.018}$ $0.949_{\pm 0.015}$ $0.955_{\pm 0.012}$ $\mathbf{0.952_{\pm 0.014}}$
ECE $0.139_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.068_{\pm 0.048}}$ $0.025_{\pm 0.009}$ $0.021_{\pm 0.006}$ $0.020_{\pm 0.009}$ $0.026_{\pm 0.010}$
LLama2-Chat-7B ACC $0.867_{\pm 0.000}$ $0.951_{\pm 0.008}$ $0.942_{\pm 0.018}$ $\mathbf{0.953_{\pm 0.012}}$ $0.952_{\pm 0.016}$ $0.952_{\pm 0.015}$
ECE $0.039_{\pm 0.000}$ $0.033_{\pm 0.006}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.044_{\pm 0.015}}$ $0.032_{\pm 0.012}$ $0.035_{\pm 0.014}$ $0.037_{\pm 0.014}$
LLama2-7B ACC $0.530_{\pm 0.000}$ $0.754_{\pm 0.140}$ $0.829_{\pm 0.121}$ $0.874_{\pm 0.105}$ $0.904_{\pm 0.062}$ $\mathbf{0.925_{\pm 0.045}}$
ECE $0.018_{\pm 0.000}$ $0.180_{\pm 0.058}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.119_{\pm 0.076}}$ $0.085_{\pm 0.072}$ $0.062_{\pm 0.027}$ $0.040_{\pm 0.012}$
Mistral-7B-v0.1 ACC $0.563_{\pm 0.000}$ $0.958_{\pm 0.007}$ $0.941_{\pm 0.058}$ $0.956_{\pm 0.030}$ $0.961_{\pm 0.022}$ $\mathbf{0.969_{\pm 0.006}}$
ECE $0.058_{\pm 0.000}$ $0.133_{\pm 0.029}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.086_{\pm 0.033}}$ $0.078_{\pm 0.032}$ $0.072_{\pm 0.032}$ $0.052_{\pm 0.020}$
vicuna-7b-v1.5 ACC $0.893_{\pm 0.000}$ $0.946_{\pm 0.011}$ $0.956_{\pm 0.008}$ $0.961_{\pm 0.012}$ $0.964_{\pm 0.008}$ $\mathbf{0.968_{\pm 0.009}}$
ECE $0.041_{\pm 0.000}$ ${\color[rgb]{0.89,0.44,0.48}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.44,0.48}0.032_{\pm 0.009}}$ $0.025_{\pm 0.008}$ $0.026_{\pm 0.006}$ $0.023_{\pm 0.005}$ $0.021_{\pm 0.005}$
Table 11: The full results (mean and standard deviation) for various experimental configurations extending Figure 3.
Dataset Case Reasoning $\hat{y}$ $p$
Strategy QA Question: Is the voice of the Genie from Disney’s Aladdin still alive? Choose the answer from True and False
Options: A) True, B) False
Ground Truth: A
7B The voice of the Genie from Disney’s Aladdin is still alive. B 0.6249
13B The voice of the Genie from Disney’s Aladdin is Robin Williams. Robin Williams died on August 11, 2014. B 0.7654
30B Robin Williams was the voice of the Genie in Disney’s Aladdin. He died on August 11, 2014. B 0.8471
World Tree
Question: A bird eats a brightly-colored insect that tastes bad.
The bird avoids preying on this type of insect again. This example is best described as ?
Options: A) a physical characteristic, B) a learned behavior, C) an inherited behavior, D) an adaptation
Ground Truth: B
7B ing the juice from fruit is a kind of physical change. A physical change is a change in the physical state of a substance. A physical change is not a change in the chemical composition of a substance. A physical change is not a change in the chemical composition of a substance. D 0.5249
13B An adaptation is a change in an organism that makes it better suited to its environment. D 0.7651
30B Adaptation means a change in the body of an organism that helps it survive in its environment. Adaptation is a kind of behavior. D 0.8764
Table 12: Qualitative Results of LLaMA on Strategy QA and WorldTree